# Certifying nonlocal properties of noisy quantum operations

### v1

@@ III.2 Local measurement protocols without entanglement
which was introduced in [8]. This can be obtained from Eq. (20) by considering a product state $\rho_{ARBS}=|00\rangle\langle 00|_{AR}\otimes|00\rangle\langle 00|_{BS}$, so that the channels $\Lambda_{AR}^{x}$ and $\Phi_{BS}^{y}$ act as state preparation channels producing input states $\rho_{AR}^{x}$ and $\sigma_{BS}^{y}$. LOSR channels produce local probability distributions from Eq. (26), as shown in [8]. This result can be extended for other classes of channels.
@@
Proof of (i). Probability distributions obtained from Eq. (26) with local operation factorize, thus LOSR channels give local probability distributions.
Proof of (ii). Similarly as in the proof of Prop. 1, define
@@
Proof of (iii). This follows from Proposition 1. ∎
Two remarks are now in order. On the one hand, note that the converse of Corollary 2 is not true. In particular, channels denoted as local-limited in [8] always produce local correlations for any choice of input states and measurements. Examples are LOSE channels where Alice and Bob share an entangled state that cannot produce nonlocal correlations [6]. On the other hand, Eq. (28) implies that a bipartite LOSE channel produces nonlocal correlations if the shared state $\tau_{EF}$ in Eq. (14) is nonlocal in some Bell protocol. This in turn dictates that if the shared state in one realization of an LOSE channel is Bell nonlocal, the channel cannot be realized by some shared local state.
III.3 A Buscemi-like characterization of LOSR operations
The results of Buscemi [15] extend to more general processes [17]. In particular, for a fixed channel ${\cal E}$ and a fixed Bell functional $\gamma$, we denote as $\gamma_{\max}({\cal E})$ the maximal value of $\gamma(p)$ that can be achieved when the conditional probability distribution $p$ is obtained by a fixed set of input states $\{\rho^{x}\}$ and $\{\sigma^{y}\}$ over all possible measurements $\{M^{a}\}$ and $\{N^{b}\}$, i.e.,
$\gamma_{\max}({\cal E}):=\max_{\{M^{a},N^{b}\}}\sum_{xy,ab}\gamma_{ab,xy}p(a,b|x,y)\,,$ (29)
where $p(ab|xy)$ is obtained through Eq. (26).
The setup depicted in Fig. 1 (b) can be seen as a particular case of so-called semi-quantum games, which are known to characterize convertibility of general processes through free operations in a generalized theory of nonlocality [17]. In particular, the setup described in Eq. (26) and Fig. 1 (b) provides semi-quantum games characterizing local transformations between LOSR quantum channels. The following follows from [17, Corollary 1]:
@@ III.3 A Buscemi-like characterization of LOSR operations
The proof of the only if direction can be easily seen as follows. If ${\cal E}$ can be converted into ${\cal E}^{\prime}$ by a product of superchannels (or a convex combination of them), then we can absorb such product of superchannels in the state preparations and measurements of semiquantum games, described in Eq. (26) and depicted in Fig. 1 (b). Then it becomes clear that $\gamma_{\max}({\cal E})\geq\gamma_{\max}({\cal E}^{\prime})$ for all functionals $\gamma$. The proof for the if direction is based on the observation that Eq. (26) gives a distribution after local superchannels are performed in the form of $\Xi_{A}\otimes\Xi_{B}[{\cal E}_{AB}]$, thus preserving the set of LOSR operations [17]. An explicit constructive proof is given in Appendix A.
Similarly as in [15], this means that a quantum channel ${\cal E}$ is not LOSR if and only if it has a strictly larger value $\gamma_{\max}({\cal E})$ than any LOSR operation for some Bell functional $\gamma$ and quantum inputs. The derivations in Appendix A provide explicit necessary and sufficient conditions to certify non-LOSR operations. Consider a set of fixed states $\{\rho^{x}_{RA_{0}}\}$ and $\{\sigma^{y}_{SB_{0}}\}$ that are linearly independent and span the whole input spaces of linear operators acting on $\mathcal{H}_{RA_{0}}$ and $\mathcal{H}_{SB_{0}}$, and assume $\mathcal{H}_{A_{0}}\approx\mathcal{H}_{R}$ and $\mathcal{H}_{B_{0}}\approx\mathcal{H}_{S}$. Let $\{\Psi^{a}_{RA_{1}}\}$ and $\{\Psi^{b}_{SB_{1}}\}$ be projective measurements on a generalized Bell basis and denote
@@
This setup allows to formulate the problem of whether or not a channel ${\cal E}_{AB}$ is LOSR as follows: A quantum channel ${\cal E}_{AB}$ is not an LOSR operation if and only if for any possible LOSR operation ${\cal E}^{\prime}_{A^{\prime}B^{\prime}}$ and for any choice of measurements $M^{a}_{RA_{1}}$ and $N^{b}_{SB_{1}}$, the distribution $p_{\cal E}(ab|xy)$ of Eq. (30) cannot be obtained via $\mathrm{Tr}(M^{a}_{RA_{1}}\otimes N^{b}_{SB_{1}}({\cal E}^{\prime}_{AB}\otimes\operatorname{id}_{RS})[\rho^{x}_{RA_{0}}\otimes\sigma^{y}_{SB_{0}}])$. By absorbing the adjoint ${{\cal E}^{\prime}_{AB}}^{\dagger}$ in the measurements $M^{a}\otimes N^{b}$, namely defining $\tilde{M}^{a}\otimes\tilde{N}^{b}:={\cal E}^{\dagger}\otimes\text{id}(M^{a}\otimes N^{b})$, one arrives at the following criterion.
@@
A bipartite quantum channel ${\cal E}_{AB}$ is not LOSR if and only for some functional $\gamma$
$\gamma(p_{\cal E})>\max_{\tilde{M}^{a},\tilde{N}^{b}}\sum_{ab,xy}\gamma_{ab,xy}\mathrm{Tr}(\tilde{M}^{a}\rho^{x})\mathrm{Tr}(\tilde{N}^{b}\sigma^{y})$ (31)
@@
This provides a separation between the set of LOSR operations and the rest, which relies only in the input/outcome probabilities and the trusted set of input states $\rho^{x}$ and $\sigma^{y}$. Therefore, it constitutes a measurement-device independent characterization of the set of LOSR operations. Note that the left hand side of Eq. (31) is given by the set of input states and the channel under consideration, and the right hand side can be upper-bounded with semidefinite programming [73].
Figure 3: Models of decoherence occurring by dephasing noise under consideration. (a) Dephasing occurs both in the inputs and the outputs as a suprechannel $\Xi_{AB}[{\cal E}_{AB}]$. Its action on the Choi matrix is given by a Gram matrix $G_{A_{0}A_{1}B_{0}B_{1}}$ without any tensor product assumption through a bipartite Eq. (17) with a bipartite structure. (b) Nonlocal dephasing occurs as independent pre- and post-processing by composition with a dephasing channel, ${\cal E}_{AB}^{\mathcal{D}}=\mathcal{D}^{G}\circ{\cal E}\circ\mathcal{D}^{G^{\prime}}$. Its action on the Choi matrix is given by a product of Gram matrix $G_{A_{0}B_{0}}\otimes G_{A_{1}B_{1}}$ through Eq. (18) with a bipartite structure. (c) Local dephasing occurs at each party independently, according to Eq. (32). (d) Local complete dephasing occurs, where the input and output states are diagonal density matrices. The last scheme representing complete decoherence corresponds to Fig 1(c), since local projective measurements in a fixed basis occur both before and after the channel is applied.
IV The role of dephasing noise
IV.1 Gradual dephasing
@@
Let us now consider the case in which the two parties $A$ and $B$ sharing a quantum state $\rho_{AB}$ or implementing a quantum channel ${\cal E}_{AB}$ belong to a joint subsystem. In this case, dephasing can occur as a joint operation (Fig. 3 (a & b)), and thus one might expect that nonlocal phenomena can occur. This can be summarized as follows.
@@
which defines a Gram matrix $G=4|\psi\rangle\langle\psi|$ with entries $G_{ij}=\left\langle v_{i}\right|v_{j}\rangle$ where $\left|v_{1}\right\rangle=\left|v_{2}\right\rangle=\left|v_{3}\right\rangle=-\left|v_{4}\right\rangle=(1,1,1,-1)^{T}/2$. Since $\left|\psi\right\rangle$ is maximally entangled, it maximally violates Clauser-Horne-Shimony-Holt (CHSH) inequality,
@@
To better understand the role of gradual decoherence aside from extremal cases, we consider two parametric types of dephasing models. On one hand, $\mathcal{D}_{q}$ damps all off-diagonal terms equally, $\mathcal{D}_{q}(\rho)=(1-q)\rho+q\operatorname{diag}(\rho)$. On the other hand, $\mathcal{D}^{\prime}_{p}$ keeps all entries except for the coherences to the fourth population level, namely $\mathcal{D}^{\prime}_{p}(\rho)_{ij}=(1-2p)\rho_{ij}$ if either $i=4\neq j$ or $j=4\neq i$.
@@
[…]
IV.2 Complete decoherence
Figure 5: Complete decoherence of states, channels and superchannels. Blue convex sets in the left represent sets of quantum states consisting of $4$, $2$ and $1$ subsystems of size $d$ related by partial trace (directed dashed arrow). Bidirectional solid arrows denote the Choi-Jamiołkowski isomorphism (C-J): a cross section of the set $\mathcal{M}_{d^{2}}$ defines through Eq. (2) the set ${\mathcal{E}}_{d}$ of quantum channels (blue section of the green convex set). An extension of the Choi-Jamiołkowski isomorphism (eC-J) links the set of two-qudit channels ${\mathcal{E}}_{d^{2}}$ with one-qudit superchannels ${\Theta}_{d}$ through Eq. (4c). Under decoherence (red thick arrows) these objects become classical polytopes depicted in red and contained inside larger orange polytopes: a quantum state $\rho\in\mathcal{M}_{d}$ is mapped to a vector $p=\operatorname{diag}(\rho)\in\Delta_{d}$, a quantum channel ${\cal E}\in\mathcal{E}_{d}$ is mapped into a stochastic matrix $S\in\mathcal{S}_{d}$, while a quantum superchannel from $\gamma_{d}$ is transformed into an element of the set $\mathcal{T}_{d}$ of classical supermaps. The sets $\mathcal{M}_{d^{2}}$ of bipartite quantum states and $\mathcal{S}_{d^{2}}$ of bipartite stochastic matrices are both obtained from the set $\mathcal{M}_{d^{4}}$ of four-partite states and inherit some of its properties.
@@ IV.2 Complete decoherence
where $(a,b,x,y)\in[d_{A_{1}}]\times[d_{B_{1}}]\times[d_{A_{0}}]\times[d_{B_{0}}]$ define the canonical basis (Fig. 5). Therefore, the decoherent action (39) defines probabilities $p(a,b|x,y)=S_{ab,xy}$ of obtaining classical outputs $ab$ conditioned to classical inputs $xy$, through the scheme of Fig. 1 (c). In  [8, Propositions 3-5] it was shown that the sets of L, Q and NS correlations correspond to the sets of correlations that can be obtained respectively from LOSR, LOSE, and qNS channels as correlations in the inputs and outputs when all measurement bases are allowed. A similar result can be formulated for a fixed basis, namely considering the decoherent action of Eq. (39), as follows.
@@
Proposition 7 can be obtained from [8, Prop. 3, 4 & 5], by absorbing the basis choice into the existence conditions. An explicit constructive proof is given in Appendix B. This provides a simple criterion to detect certain classes of quantum channels, from local measurements in the computational basis: One certifies that a quantum channel ${\cal E}_{AB}$ is outside of the set of LOSR (LOSE, qNS) channels if its decoherent action $S_{AB}^{\mathcal{E}}$ is outside of the set of L (Q, NS) correlations. This criterion can be implemented in decoherent setups, as it involves only measurements in the computational basis.
Let us highlight that the converse is not true: there exist non-LOSR quantum channels with local decoherent action. As an example, consider the 2-qubit channel with unitary (single) Kraus operator $U=\sum_{i,j=0}^{1}\left|\phi_{ij}\right\rangle\left\langle i+j,j\right|$, which maps the computational basis to the Bell states $\left|\phi_{ij}\right\rangle=\sigma_{X}^{i}\sigma_{Z}^{j}\otimes\mathbb{1}_{2}(\left|00\right\rangle+\left|11\right\rangle)/\sqrt{2}$ where $\sigma_{X}$ and $\sigma_{Z}$ are Pauli matrices. Even though $U$ can generate entanglement, its decoherent action $S^{U}=U\odot U^{*}=\frac{1}{2}\left(\mathbb{1}_{2}^{\otimes 2}+\sigma_{X}^{\otimes 2}\right)$ is local. However, note that this criterion can be enhanced with any operations that preserve the set of LOSR. In the example above, the freedom of local measurement basis considered in [8] is enough: the bipartite unitary $U$ can be transformed by local unitaries to $\tilde{U}=(H\otimes H)U(\mathbf{1}\otimes H)$, where $H$ is the Hadamard gate with $H\left|0\right\rangle=\left|+\right\rangle$ and $H\left|1\right\rangle=\left|-\right\rangle$. We have $S^{\tilde{U}}=\tilde{U}\odot\tilde{U}^{*}=|0\rangle\langle 0|+|1\rangle\langle 3|+|2\rangle\langle 2|+|3\rangle\langle 1|$ which defines extremal nonsignaling correlations.
@@
By Proposition 7, Bell functionals on the decoherent action of a quantum channel can be used to construct linear nonlocality witnesses, similar to entanglement witnesses [34]. This mechanism is illustrated in Fig. 2. Let $\mathcal{L}$ be one of the three subsets of bipartite quantum channels: LOSR, LOSE, and qNS. Let $S$ be the corresponding set of bipartite probability distributions under the correspondence in Prop. 7, namely, L, Q, and NS. Consider
@@
For the CHSH functional, $\gamma_{ab,xy}=(-1)^{a+b}(-1)^{xy}$, it holds that $\gamma_{\operatorname{L{}}}=2$, $\gamma_{\operatorname{Q}}=2\sqrt{2}$ and $\gamma_{\operatorname{NS}}=4$.
@@
On the other hand, by the isomorphism of Proposition 7, any bipartite stochastic matrix in a distinguished subset of correlations can be obtained from decoherence of a quantum channel in a corresponding subset. An explicit example is the convex hull of bipartite stochastic matrices (namely, bipartite probability distributions) $pR+(1-p)S$ with
$R=\frac{1}{2}\begin{pmatrix}1&1&1&0\\ 0&0&0&1\\ 0&0&0&1\\ 1&1&1&0\end{pmatrix},\,\,S=\frac{1}{2}\begin{pmatrix}0&0&0&1\\ 1&1&1&0\\ 1&1&1&0\\ 0&0&0&1\\ \end{pmatrix},$ (43)
which attains the maximal quantum value $\gamma_{Q}=2\sqrt{2}$ of CHSH for $p=(1+1/\sqrt{2})/2$.
Figure 6: Partial cross section of the space of bipartite distributions spanned by $\mathbf{1}_{4}$, $R$ and $S$ in Eq. (43), $P(s,t)=sR+tS+(1-s-t)\mathbf{1}_{4}$ where $s+t\leq 1$. The distribution $\mathbf{1}_{4}$ is local, while $R$ and $S$ are extremal nonsignaling (PR-boxes). The orange region corresponds with distributions that are local, while dark and light purple regions denote first (NPA1) and second (NPA2) level of Navascués-Pironio-Acín (NPA) approximation [77] to the set $Q$ of quantum distributions (not plotted). Green region depicts the remaining non-signaling distributions. Each of these subsets of distributions can be obtained by LOSR, LOSE and qNS channels under complete decoherence.
@@
In Fig. 6 we show a two dimensional slice of the polytope of distributions consisting of the convex hull of $R$, $S$ and $\mathbf{1}$ containing different subsets. The sets of local (L) and nonsignaling (NS) distributions are bounded by linear programming, while the set of quantum (Q) distributions is approximated by the first two levels of the Navascués-Pironio-Acín (NPA) hierarchy [77]. The subsets L, Q and NS of distributions are obtained by decoherence of the subsets LOSR, LOSE and qNS of channels.
V Conclusions and open questions

### v3

@@ 3.2 Local measurement protocols without entanglement
which was introduced in [7]. This can be obtained from Eq. (20) by considering a product state $\rho_{ARBS}=|00\rangle\langle 00|_{AR}\otimes|00\rangle\langle 00|_{BS}$, so that the channels $\Lambda_{AR}^{x}$ and $\Phi_{BS}^{y}$ act as state preparation channels producing input states $\rho_{AR}^{x}$ and $\sigma_{BS}^{y}$. LOSR channels produce local probability distributions from Eq. (26), as shown in [7]. In items $(ii)$ and $(iii)$ of the following Corollary we extend this result to other classes of channels.
@@
Proof of (i). Notice that probability distributions in Eq. (26) from LOSR channels give local probability distributions. This was noted earlier in [7, see Fig. 8].
Proof of (ii). Similarly as in Prop. 1, define
@@
Item (iii) follows from Proposition 1. ∎
Two remarks are now given. On the one hand, note that the converse of Corollary 2 is not true. In particular, channels denoted as local-limited in [7] always produce local correlations for any choice of input states and measurements. Examples are LOSE channels where Alice and Bob share an entangled state that cannot produce nonlocal correlations [5]. On the other hand, Eq. (28) implies that a bipartite LOSE channel produces nonlocal correlations if the shared state $\tau_{EF}$ in Eq. (14) is nonlocal in some Bell protocol. This in turn dictates that if the shared state in one realization of an LOSE channel is Bell nonlocal, the channel cannot be realized by some shared local state.
3.3 A Buscemi-like characterization of LOSR operations
The results of Buscemi [14] extend to more general processes [16]. For a fixed channel ${\cal E}$ and Bell functional $\gamma$, we denote as $\gamma_{\max}({\cal E})$ the maximal value of $\gamma(p)$ in Eq. (10) that can be achieved when the conditional probability distribution $p$ is obtained through Eq. (26) by a fixed set of input states $\rho^{x}$ and $\sigma^{y}$ over all possible measurements $M^{a}$ and $N^{b}$. That is,
$\gamma_{\max}({\cal E}):=\max_{\{M^{a},N^{b}\}}\sum_{xy,ab}\gamma_{ab,xy}p(a,b|x,y)\,.$ (29)
The setup depicted in Fig. 1 (b) can be seen as a particular case of so-called semi-quantum games, which are known to characterize convertibility of general processes through free operations in a generalized theory of nonlocality [16]. In particular, the setup described in Eq. (26) and Fig. 1 (b) provides semi-quantum games characterizing local transformations between LOSR quantum channels. The following is a particular case of [16, Corollary 1]:
@@ 3.3 A Buscemi-like characterization of LOSR operations
Proof.
Starting from the more general results in [16], the proof of the only if direction can be easily seen as follows. If ${\cal E}$ can be converted into ${\cal E}^{\prime}$ by a product of superchannels (or a convex combination of them), then we can absorb such product of superchannels in the state preparations and measurements of semiquantum games, described in Eq. (26) and depicted in Fig. 1 (b). Then it becomes clear that $\gamma_{\max}({\cal E})\geq\gamma_{\max}({\cal E}^{\prime})$ for all functionals $\gamma$. The proof for the if direction is based on the observation that Eq. (26) gives a distribution after local superchannels are performed in the form of $\Xi_{A}\otimes\Xi_{B}[{\cal E}_{AB}]$, thus preserving the set of LOSR operations [16]. For self-consistency of the paper, an explicit constructive proof is given in Appendix A. ∎
Similarly as in [14], this means that a quantum channel ${\cal E}$ is not LOSR if and only if for some Bell functional $\gamma$ and quantum inputs, the value $\gamma_{\max}({\cal E})$ in Eq. (29) cannot be attained with LOSR operations. The derivations in Appendix A provide explicit necessary and sufficient conditions to certify non-LOSR operations. Consider a set of fixed states $\{\rho^{x}_{RA_{0}}\}$ and $\{\sigma^{y}_{SB_{0}}\}$ that are linearly independent and span the whole input spaces of linear operators acting on $\mathcal{H}_{RA_{0}}$ and $\mathcal{H}_{SB_{0}}$, and assume $\mathcal{H}_{A_{0}}\approx\mathcal{H}_{R}$ and $\mathcal{H}_{B_{0}}\approx\mathcal{H}_{S}$. Let $\{\Psi^{a}_{RA_{1}}\}$ and $\{\Psi^{b}_{SB_{1}}\}$ be projective measurements on a generalized Bell basis and denote
@@
This setup allows to formulate the problem of whether or not a channel ${\cal E}_{AB}$ is LOSR as follows: A quantum channel ${\cal E}_{AB}$ is not an LOSR operation if and only if for any possible LOSR operation ${\cal E}^{\prime}_{A^{\prime}B^{\prime}}$ and for any choice of measurements $M^{a}_{RA_{1}}$ and $N^{b}_{SB_{1}}$, the distribution $p_{\cal E}(ab|xy)$ of Eq. (30) cannot be obtained via $\mathrm{Tr}(M^{a}_{RA_{1}}\otimes N^{b}_{SB_{1}}({\cal E}^{\prime}_{AB}\otimes\operatorname{id}_{RS})[\rho^{x}_{RA_{0}}\otimes\sigma^{y}_{SB_{0}}])$. By absorbing the adjoint ${{\cal E}^{\prime}_{AB}}^{\dagger}$ in the measurements $M^{a}\otimes N^{b}$, namely defining $\tilde{M}^{a}\otimes\tilde{N}^{b}:={\cal E}^{\dagger}\otimes\text{id}(M^{a}\otimes N^{b})$, we derive the following criterion, analogous to Buscemi’s characterization of entangled states  [14].
@@
A bipartite quantum channel ${\cal E}_{AB}$ is not LOSR if and only if for some linear functional $\gamma$,
$\gamma(p_{\cal E})>\max_{\tilde{M}^{a},\tilde{N}^{b}}\sum_{ab,xy}\gamma_{ab,xy}\mathrm{Tr}(\tilde{M}^{a}\rho^{x})\mathrm{Tr}(\tilde{N}^{b}\sigma^{y})\,,$ (31)
@@
This provides a separation between the set of LOSR operations and the rest, which relies only in the input/outcome probabilities and the trusted set of input states $\rho^{x}$ and $\sigma^{y}$. Therefore, it constitutes a measurement-device-independent characterization of the set of LOSR operations. Note that the left hand side of Eq. (31) is given by the set of input states and the channel under consideration, and the right hand side can be upper-bounded with semidefinite programming [75].
Figure 3: Models of decoherence occurring by dephasing noise under consideration. (a) Dephasing occurs both in the inputs and the outputs as a superchannel $\Xi_{AB}[{\cal E}_{AB}]$. Its action on the Choi matrix is given by a Gram matrix $G_{A_{0}A_{1}B_{0}B_{1}}$ without any tensor product assumption through a bipartite Eq. (17) with a bipartite structure. (b) Nonlocal dephasing occurs as independent pre- and post-processing by composition with a dephasing channel, ${\cal E}_{AB}^{\mathcal{D}}=\mathcal{D}^{G}\circ{\cal E}\circ\mathcal{D}^{G^{\prime}}$. Its action on the Choi matrix is given by a product of Gram matrix $G_{A_{0}B_{0}}\otimes G_{A_{1}B_{1}}$ through Eq. (18) with a bipartite structure. (c) Local dephasing occurs at each party independently, according to Eq. (32). (d) Local complete dephasing occurs, where the input and output states are diagonal density matrices. The last scheme representing complete decoherence corresponds to Fig. 1 (c), since local projective measurements in a fixed basis occur both before and after the channel is applied.
4 The role of dephasing noise
4.1 Gradual dephasing
@@
Let us now consider the case in which the two parties $A$ and $B$ sharing a quantum state $\rho_{AB}$ or implementing a quantum channel ${\cal E}_{AB}$ belong to a joint subsystem. In this case, dephasing can occur as a joint operation (Fig. 3 (a & b)), and thus one might expect that nonlocal phenomena can occur. This leads us to the following result.
@@
which defines a Gram matrix $G=4|\psi\rangle\langle\psi|$ with entries $G_{ij}=\left\langle v_{i}\right|v_{j}\rangle$ where $\left|v_{1}\right\rangle=\left|v_{2}\right\rangle=\left|v_{3}\right\rangle=-\left|v_{4}\right\rangle=(1,1,1,-1)^{T}/2$. Since $\left|\psi\right\rangle$ is maximally entangled, it maximally violates CHSH inequality,
@@
To better understand the role of gradual decoherence aside from extremal cases, we consider two parametric types of dephasing models. On the one hand, $\mathcal{D}_{q}$ damps all off-diagonal terms equally, $\mathcal{D}_{q}(\rho)=(1-q)\rho+q\operatorname{diag}(\rho)$. On the other hand, $\mathcal{D}^{\prime}_{p}$ keeps all entries except for the coherences to the fourth population level, namely $\mathcal{D}^{\prime}_{p}(\rho)_{ij}=(1-2p)\rho_{ij}$ if either $i=4\neq j$ or $j=4\neq i$.
@@
Figure 4: Entanglement generated by dephasing. The parameters $p$ and $q$ quantify the strength of the dephasing channels $\mathcal{D}_{q}$ and $\mathcal{D}^{\prime}_{p}$ acting on the maximally coherent state $\rho=|++\rangle\langle++|$ [76, 52], which can be mapped to less coherent but entangled states. The color bar evaluates the negativity of the partial transpose of the resulting state, indicating maximal entanglement when taking the value $0.5$.
4.2 Complete decoherence
Figure 5: Complete decoherence of states, channels and superchannels. Blue convex sets in the left represent sets of quantum states consisting of $4$, $2$ and $1$ subsystems of size $d$ related by partial trace (directed dashed arrow). Bidirectional solid arrows denote the Choi-Jamiołkowski isomorphism (C-J): a cross section of the set $\mathcal{M}_{d^{2}}$ defines through Eq. (2) the set ${\mathcal{E}}_{d}$ of quantum channels (blue section of the green convex set). An extension of the Choi-Jamiołkowski isomorphism (eC-J) links the set of two-qudit channels ${\mathcal{E}}_{d^{2}}$ with one-qudit superchannels ${\Theta}_{d}$ through Eq. (4c). Under decoherence (red thick arrows) these objects become classical polytopes depicted in red and contained inside larger orange polytopes: a quantum state $\rho\in\mathcal{M}_{d}$ is mapped to a vector $p=\operatorname{diag}(\rho)\in\Delta_{d}$, a quantum channel ${\cal E}\in\mathcal{E}_{d}$ is mapped into a stochastic matrix $S\in\mathcal{S}_{d}$, while a quantum superchannel from $\Theta_{d}$ is transformed into an element of the set $\mathcal{T}_{d}$ of classical supermaps. The sets $\mathcal{M}_{d^{2}}$ of bipartite quantum states and $\mathcal{S}_{d^{2}}$ of bipartite stochastic matrices are both obtained from the set $\mathcal{M}_{d^{4}}$ of four-partite states and inherit some of its properties.
@@ 4.2 Complete decoherence
where $(a,b,x,y)\in[d_{A_{1}}]\times[d_{B_{1}}]\times[d_{A_{0}}]\times[d_{B_{0}}]$ define the canonical basis (Fig. 5). Therefore, the decoherent action (39) defines probabilities $p(a,b|x,y)=S_{ab,xy}$ of obtaining classical outputs $ab$ conditioned to classical inputs $xy$, through the scheme of Fig. 1 (c). In  [7, Propositions 3-5] it was shown that the sets of L, Q and NS correlations correspond to the sets of correlations that can be obtained respectively from LOSR, LOSE, and QNS channels as correlations in the inputs and outputs when all measurement bases are allowed. Since we are interested in dephasing noise, here we reformulate this result for a fixed basis considering the decoherent action of Eq. (39), as follows.
@@
Proposition 7 can be obtained from [7, Prop. 3, 4 & 5], by absorbing the basis choice into the existence conditions. A qualitative analysis of the channel nonlocality that is witnessed by the nonlocality of such correlations is carried out in a different work [77, Prop. 1]. For self-consistency, an explicit constructive proof is given in Appendix B. This provides a simple criterion to detect certain classes of quantum channels, from local measurements in the computational basis: one certifies that a quantum channel ${\cal E}_{AB}$ is outside of the set of LOSR (LOSE, QNS) channels if its decoherent action $S_{AB}^{\mathcal{E}}$ is outside of the set of L (Q, NS) correlations. This criterion can be implemented in decoherent setups, as it involves only measurements in the computational basis.
Let us remark that the converse is not true: there exist non-LOSR quantum channels with local decoherent action. As an example, consider the 2-qubit channel with unitary (single) Kraus operator $U=\sum_{i,j=0}^{1}\left|\phi_{ij}\right\rangle\left\langle i+j,j\right|$, which maps the computational basis to the Bell states $\left|\phi_{ij}\right\rangle=\sigma_{X}^{i}\sigma_{Z}^{j}\otimes\mathbf{1}(\left|00\right\rangle+\left|11\right\rangle)/\sqrt{2}$ where $\sigma_{X}$ and $\sigma_{Z}$ are Pauli matrices. Even though $U$ can generate entanglement, its decoherent action $S^{U}=U\odot U^{*}=\frac{1}{2}\left(\mathbf{1}^{\otimes 2}+\sigma_{X}^{\otimes 2}\right)$ is local. However, note that this criterion can be enhanced with any operations that preserve the set of LOSR. In the example above, the freedom of local measurement basis considered in [7] is enough: the bipartite unitary $U$ can be transformed by local unitaries to $\tilde{U}=(H\otimes H)U(\mathbf{1}\otimes H)$, where $H$ is the Hadamard gate with $H\left|0\right\rangle=\left|+\right\rangle$ and $H\left|1\right\rangle=\left|-\right\rangle$. We have $S^{\tilde{U}}=\tilde{U}\odot\tilde{U}^{*}=|0\rangle\langle 0|+|1\rangle\langle 3|+|2\rangle\langle 2|+|3\rangle\langle 1|$ which defines extremal nonsignaling correlations.
@@
By Proposition 7, Bell functionals on the decoherent action of a quantum channel can be used to construct linear nonlocality witnesses, similar to entanglement witnesses [36]. This mechanism is illustrated in Fig. 2. Let $\mathcal{L}$ be one of the three subsets of bipartite quantum channels: LOSR, LOSE, and QNS. Let $S$ be the corresponding set of bipartite probability distributions under the correspondence in Proposition 7, namely, L, Q, and NS. Consider
@@
For the CHSH functional, $\gamma_{ab,xy}=(-1)^{a+b}(-1)^{xy}$, it holds that $\gamma_{\operatorname{L{}}}=2$, $\gamma_{\operatorname{Q}}=2\sqrt{2}$ and $\gamma_{\operatorname{NS}}=4$ [34, 13].
@@
By the isomorphism of Proposition 7, any bipartite stochastic matrix in a distinguished subset of correlations can be obtained from decoherence of a quantum channel in a corresponding subset. An explicit example is the convex hull of bipartite stochastic matrices (namely, bipartite probability distributions) $pR+(1-p)S$ with
$R=\frac{1}{2}\begin{pmatrix}1&1&1&0\\ 0&0&0&1\\ 0&0&0&1\\ 1&1&1&0\end{pmatrix},\,S=\frac{1}{2}\begin{pmatrix}0&0&0&1\\ 1&1&1&0\\ 1&1&1&0\\ 0&0&0&1\\ \end{pmatrix}$ (43)
@@
which attains the maximal quantum value $\gamma_{Q}=2\sqrt{2}$ of CHSH for $p=(1+1/\sqrt{2})/2$. In Fig. 6 we show a two-dimensional slice of the polytope of distributions consisting of the convex hull of $R$, $S$ and $\mathbf{1}$ containing different subsets. The sets of local (L) and nonsignaling (NS) distributions are bounded by linear programming, while the set of quantum (Q) distributions is approximated by the first two levels of the Navascués-Pironio-Acín (NPA) hierarchy [80]. The subsets L, Q and NS of distributions are obtained by decoherence of the subsets LOSR, LOSE and QNS of channels.
Figure 6: Partial cross section of the space of bipartite distributions spanned by $\mathbf{1}_{4}$, $R$ and $S$ in Eq. (43), $P(s,t)=sR+tS+(1-s-t)\mathbf{1}_{4}$ where $s+t\leq 1$. The distribution $\mathbf{1}_{4}$ is local, while $R$ and $S$ are Popescu-Rohrlich (PR) correlations [87], which are extremal nonsignaling. The orange region corresponds with distributions that are local, while dark and light purple regions denote first (NPA1) and second (NPA2) level of Navascués-Pironio-Acín (NPA) approximation [80] to the set $Q$ of quantum distributions (not plotted). Green region depicts the remaining nonsignaling distributions. Each of these subsets of distributions can be obtained by LOSR, LOSE and QNS channels under complete decoherence.
5 Conclusions and open questions